# biblio.bib

### v2

@@ 5 Mapping class groups of trees
Let $T$ be a locally finite infinite tree. Recall that the end space $E\coloneqq\partial T$ is homeomorphic to a closed subset of the Cantor set. Moreover, if we equip $\operatorname{Homeo}(E)$ with compact-open topology, then [bestvina2025groups, Corollary 4.12] states that the topological groups $\operatorname{\mathcal{MCG}}(T)\simeq\operatorname{Homeo}(E)$ are isomorphic.
Hence, studying the amenability of the mapping class group of a tree $T$ is equivalent to understanding the same problem for the group of homeomorphisms of a closed subset $E$ of the Cantor set. In view of this, it is convenient to appeal to Cantor-Bendixson analysis, for which more details can be found in [kechris2012classical, §I.6].
@@
$X^{(\alpha)}$ $=\bigcap_{\beta<\alpha}X^{(\beta)},\text{ if }\alpha\text{ is limit.}$.
@@
For each countable ordinal $\alpha\geq 0$, we define $G_{\alpha}<\operatorname{Homeo}(E)\simeq\operatorname{\mathcal{MCG}}(T)$ to be the pointwise stabiliser of $E^{(\alpha)}$. In particular, we note that if $\alpha<\beta$, then $G_{\alpha}\subset G_{\beta}$. Equivalently, $G_{\alpha}$ consists of homeomorphisms of $E$ supported on $\bigcup_{\lambda<\alpha}D_{\lambda}$.
Recall that the compact-open topology on $\operatorname{Homeo}(E)$ is such that a basis of neighbourhoods of the identity is the collection $\big{(}V_{\mathcal{P}}\big{)}_{\mathcal{P}}$, where $\mathcal{P}$ runs over the finite partition $\mathcal{P}$ of $E$ into clopen subsets and $V_{\mathcal{P}}$ os the clopen subgroup of elements leaving every $P_{i}\in\mathcal{P}$ invariant. See for example [bestvina2025groups, §4.1].
@@
First, we need to note that the closed subgroup $G_{\alpha}$ is a normal subgroup of $G_{\alpha+1}$ so that $G_{\alpha+1}/G_{\alpha}$ is a topological group. Indeed, as a pointwise stabiliser of a closed subset, $G_{\alpha}$ is defined by a closed condition and is a closed subgroup. Moreover, given any $g\in G_{\alpha}$, any $f\in G_{\alpha+1}$, and any $x\in E^{(\alpha)}$, since $f^{-1}(x)\in E^{(\alpha)}$ as we have remarked earlier, we can conclude that $fgf^{-1}(x)=f\circ g\big{(}f^{-1}(x)\big{)}=f\circ f^{-1}(x)=x$, i.e. we have $fgf^{-1}\in G_{\alpha}$.
@@
It is not difficult to see that the restriction map $\rho_{\alpha}$ is continuous, which further implies that the quotient map $\phi_{\alpha}$ is also continuous. To show that $\phi_{\alpha}$ is a homeomorphism, it suffices to show that $\phi_{\alpha}$ is also open. Take a finite partition $\mathcal{P}\coloneqq(P_{i})$ of $E$ into clopen subsets. Then $\mathcal{P}_{\alpha}\coloneqq(P_{i}\cap E^{(\alpha)})_{i}$ will become a finite partition of $E^{(\alpha)}$ into clopen subsets. By the definition of quotient topology, a basis of neighbourhoods in $\operatorname{Homeo}(E)/G_{\alpha}$ of the identity is the subgroup of classes $[\widehat{f}]$ such that $\widehat{f}\in V_{\mathcal{P}}$. Let $f\coloneqq\phi_{\alpha}([\widehat{f}])\in\operatorname{Homeo}(E^{(\alpha)})$. Then $f=\rho_{\alpha}(\widehat{f})$, which implies that
@@
Let us treat the case when $\alpha=1$.
@@
We remark that $G_{1}$ is the subgroup of homeomorphisms of $E$ that are only supported on the discrete subspace of isolated points $D_{0}$. If $|D_{0}|<\infty$, then $G_{1}$ is finite and thus is amenable. Suppose now that $|D_{0}|=\infty$. Let $F\subset D_{0}$ be an arbitrary finite subset and let $\mathfrak{S}_{F}$ be the symmetry group of $F$. We note that $\mathfrak{S}_{F}<G_{1}$ be definition and that $\mathfrak{S}_{F}$ is amenable as a finite subgroup. Now consider the subgroup
@@
We note that the subspace topology on $D_{\alpha}$ is discrete, otherwise there exists a point in $D_{\alpha}$ that is accumulation point of $D_{\alpha}$ itself, which contradicts to the definition of Cantor-Bendixson rank. So for each $p\neq q\in D_{\alpha}$, one can find a clopen neighbourhood $U\ni p$ and $V\ni q$ in $E$ so that $U\cap V=\emptyset$ and $U\simeq V\simeq\omega^{\alpha}+1$. In turn, we can construct a homeomorphism $f\colon E\to E$ such that $f(V)=U$ with $f(p)=f(q)$ and $f(x)=x$ for every $x\notin U\cup V\,$. Hence, we can conclude that the support
@@
is the full set of isolated points in $E^{(\alpha)}$, i.e. the full set of $D_{\alpha}\,$. Then apply the same arguments in Lemma 5.2 to $E^{(\alpha)}$ and $G^{\sharp}_{\alpha}$ will give the amenability of $G^{\sharp}_{\alpha}\,$. ∎
@@
Suppose that $E$ is countable. This is equivalent to $C=\emptyset$, or $E=D\,$. Then the Cantor-Bendixson decomposition yields $D=\bigcup_{0\leq\lambda\leq\alpha_{0}}D_{\lambda}$ for some countable ordinal $\alpha_{0}>0$. By definition, it immediately follows that $\operatorname{Homeo}(E)=G_{\alpha}$ for any $\alpha>\alpha_{0}\,$. Then Proposition 5.6 implies that $\operatorname{Homeo}(E)$ is amenable, or equivalent $\operatorname{\mathcal{MCG}}(T)$ of a locally finite infinite tree with countable end space $\partial T$ is amenable.
@@
By our assumption, the subset $A\coloneqq C\setminus\overline{D}\subset C$ is a non-empty open subset of the Cantor set $C$. By a classical result [schoenfeld1975alternate], this subset $A$ is either homeomorphic to $C$ itself or homeomorphic to $C\setminus\{x\}$ for any $x\in C\,$. In either case, there exists a clopen subset $K\subset A$ that is homeomorphic to the Cantor set. Moreover, by extending the homeomorphisms on $K$ by identity on $E\setminus K\,$, we can show that the action of the open subgroup $\operatorname{Homeo}(E)_{K}<\operatorname{Homeo}(E)$ on $K$ factors through $\operatorname{Homeo}(K)$ via restriction. Since $K$ is homeomorphic to the Cantor set, it follows from our discussion above that $\operatorname{Homeo}(K)\,$, and thus $\operatorname{Homeo}(E)_{K}$ admit no invariant probability measure on the compact set $K$. This implies that $\operatorname{Homeo}(E)_{K}$ is not amenable, forcing $\operatorname{Homeo}(E)$ to be non-amenable by (A1). ∎
It is worth remarking that Proposition 5.8 is not a characterisation of non-amenable cases, since there exists uncountable closed subset $E$ of the Cantor set with Cantor-Bendixson decomposition $E=C\cup D$ such that $\overline{D}\cap C=C$, yet $\operatorname{Homeo}(E)$ is not amenable. It suffices to consider the compact space $E$ of the union of the Cantor set $C$ and a countable set of isolated points $D$ such that each point of the Cantor set is an accumulation point of the isolated points. Then $\overline{D}\cap C=C$ but $\operatorname{Homeo}(E)$ is not amenable as it admits no invariant probability measure on $C$.
Using the partial order on the end space $E$ from [mann2023large], for a maximal end $x\in E$, we write $E(x)$ for the ends that are equivalent to $x$. If there exists a maximal end $x\in E$ such that $x$ is stable and $E(x)$ is homeomorphic to the Cantor set, then $\operatorname{Homeo}(E)$ acts on the clopen subset $E(x)\subset E$ transitively (see for example [mann2024two]), and thus does not have an invariant probability measure on $E(x)$, which implies the non-amenability of $\operatorname{Homeo}(E)$ by Theorem 1.4. Hence, by the dichotomy of maximal ends established in [mann2023large, Proposition 4.7], we can further conclude:
@@
[…]
Yusen Long
Université Paris-Est Créteil, CNRS, LAMA UMR8050, F-94010 Créteil, France
Email address: yusen.long@u-pec.fr

### v3

@@ 5 Mapping class groups of trees
Let $T$ be a locally finite infinite tree. Recall that the end space $E\coloneqq\partial T$ is homeomorphic to a closed subset of the Cantor set. Moreover, if we equip $\operatorname{Homeo}(E)$ with compact-open topology, then [AB25, Corollary 4.12] states that the topological groups $\operatorname{\mathcal{MCG}}(T)\simeq\operatorname{Homeo}(E)$ are isomorphic.
Hence, studying the amenability of the mapping class group of a tree $T$ is equivalent to understanding the same problem for the group of homeomorphisms of a closed subset $E$ of the Cantor set. In view of this, it is convenient to appeal to Cantor-Bendixson analysis, for which more details can be found in [KEC12, §I.6].
@@
$X^{(\alpha)}$ $=\bigcap_{\beta<\alpha}X^{(\beta)},\text{ if }\alpha\text{ is limit.}$
@@
For each countable ordinal $\alpha\geq 0$, we define $G_{\alpha}<\operatorname{Homeo}(E)\simeq\operatorname{\mathcal{MCG}}(T)$ as the pointwise stabiliser of $E^{(\alpha)}$. In particular, we note that if $\alpha<\beta$, then $G_{\alpha}\subset G_{\beta}$. Equivalently, $G_{\alpha}$ consists of homeomorphisms of $E$ supported on $\bigcup_{\lambda<\alpha}D_{\lambda}$.
Recall that the compact-open topology on $\operatorname{Homeo}(E)$ is such that a basis of neighbourhoods of the identity is the collection $\big(V_{\mathcal{P}}\big)_{\mathcal{P}}$, where $\mathcal{P}$ runs over the finite partition $\mathcal{P}$ of $E$ into clopen subsets and $V_{\mathcal{P}}$ os the clopen subgroup of elements leaving every $P_{i}\in\mathcal{P}$ invariant. See for example [AB25, §4.1].
@@
First, we need to note that the closed subgroup $G_{\alpha}$ is a normal subgroup of $G_{\alpha+1}$ so that $G_{\alpha+1}/G_{\alpha}$ is a topological group. Indeed, as a pointwise stabiliser of a closed subset, $G_{\alpha}$ is defined by a closed condition and is a closed subgroup. Moreover, given any $g\in G_{\alpha}$, any $f\in G_{\alpha+1}$, and any $x\in E^{(\alpha)}$, since $f^{-1}(x)\in E^{(\alpha)}$ by the definition of $G_{\alpha+1}$, we can conclude that $fgf^{-1}(x)=f\circ g\big(f^{-1}(x)\big)=f\circ f^{-1}(x)=x$, i.e. we have $fgf^{-1}\in G_{\alpha}$.
@@
[…]
@@
Let us treat the case where $\alpha=1$.
@@
We remark that $G_{1}$ is the subgroup of homeomorphisms of $E$ that are only supported on the discrete subspace of isolated points $D_{0}$. If $|D_{0}|<\infty$, then $G_{1}$ is finite and thus is amenable. Suppose now that $|D_{0}|=\infty$. Let $F\subset D_{0}$ be an arbitrary finite subset and let $\mathfrak{S}_{F}$ be the symmetry group of $F$. We note that $\mathfrak{S}_{F}<G_{1}$ by definition and that $\mathfrak{S}_{F}$ is amenable as a finite subgroup. Now consider the subgroup
@@
We note that the subspace topology on $D_{\alpha}$ is discrete, otherwise there exists a point in $D_{\alpha}$ that is an accumulation point of $D_{\alpha}$ itself, which contradicts the definition of Cantor-Bendixson rank. So for each $p\neq q\in D_{\alpha}$, one can find a clopen neighbourhood $U\ni p$ and $V\ni q$ in $E$ so that $U\cap V=\emptyset$ and $U\simeq V\simeq\omega^{\alpha}+1$. In turn, we can construct a homeomorphism $f\colon E\to E$ such that $f(V)=U$ with $f(p)=f(q)$ and $f(x)=x$ for every $x\notin U\cup V\,$. Hence, we can conclude that the support
@@
is the full set of isolated points in $E^{(\alpha)}$, i.e. the full set of $D_{\alpha}\,$. Then applying the same arguments in Lemma 5.2 to $E^{(\alpha)}$ and $G^{\sharp}_{\alpha}$ will give the amenability of $G^{\sharp}_{\alpha}\,$. ∎
@@
Suppose that $E$ is countable. This is equivalent to $C=\emptyset$, or E=DE=D\,. Then the Cantor-Bendixson decomposition yields $D=\bigcup_{0\leq\lambda\leq\alpha_{0}}D_{\lambda}$ for some countable ordinal $\alpha_{0}>0$. By definition, it immediately follows that $\operatorname{Homeo}(E)=G_{\alpha}$ for any $\alpha>\alpha_{0}\,$. Then Proposition 5.6 implies that $\operatorname{Homeo}(E)$ is amenable, or equivalent $\operatorname{\mathcal{MCG}}(T)$ of a locally finite infinite tree with countable end space $\partial T$ is amenable.
@@
By our assumption, the subset $A\coloneqq C\setminus\overline{D}\subset C$ is a non-empty open subset of the Cantor set $C$. By a classical result [GS75], this subset $A$ is either homeomorphic to $C$ itself or homeomorphic to $C\setminus\{x\}$ for any $x\in C\,$. In either case, there exists a clopen subset $K\subset A$ that is homeomorphic to the Cantor set. Moreover, by extending the homeomorphisms on $K$ by identity on $E\setminus K\,$, we can show that the action of $\operatorname{Homeo}(E)$ on $K$ maps surjectively via restriction to the actions of the open subgroup $\operatorname{Homeo}(E)_{K}$ on $K$. Since $K$ is homeomorphic to the Cantor set, it follows from our discussion above that $\operatorname{Homeo}(K)\,$, and thus $\operatorname{Homeo}(E)_{K}$ does not admit an invariant probability measure on the compact set $K$. This implies that $\operatorname{Homeo}(E)_{K}$ is not amenable, forcing $\operatorname{Homeo}(E)$ to be non-amenable by (A1). ∎
It is worth remarking that Proposition 5.8 is not a characterisation of non-amenable cases, since there exists an uncountable closed subset $E$ of the Cantor set with Cantor-Bendixson decomposition $E=C\cup D$ such that $\overline{D}\cap C=C$, yet $\operatorname{Homeo}(E)$ is not amenable. It suffices to consider the compact space $E$ of the union of the Cantor set $C$ and a countable set of isolated points $D$ such that each point of the Cantor set is an accumulation point of the isolated points. Then $\overline{D}\cap C=C$ but $\operatorname{Homeo}(E)$ is not amenable as it does not admit an invariant probability measure on $C$.
Using the partial order on the end space $E$ from [MR23], for a maximal end $x\in E$, we write $E(x)$ for the ends that are equivalent to $x$. If there exists a maximal end $x\in E$ such that $x$ is stable and $E(x)$ is homeomorphic to the Cantor set, then $\operatorname{Homeo}(E)$ acts on the clopen subset $E(x)\subset E$ transitively (see for example [MR24]), and thus does not have an invariant probability measure on $E(x)$, which implies the non-amenability of $\operatorname{Homeo}(E)$ by Theorem 1.6. Hence, by the dichotomy of maximal ends established in [MR23, Proposition 4.7], we can further conclude: